# GeneMamba: Efficient and Effective Foundation Model on Single Cell Data

### v1

@@ Abstract
Transformer-based architectures have shown promise across various domains but struggle with computational inefficiency and scalability. To address these challenges, we introduce GeneMamba, a novel model designed specifically for single-cell data analysis. GeneMamba incorporates the BiMamba module to efficiently capture gene context information and employs biologically meaningful loss functions during training. The model enables scalable processing of over 50 million cells while significantly reducing computational costs. It delivers strong performance in multi-batch integration, cell type annotation, and gene pair correlation analysis. Furthermore, reconstruction experiments highlight GeneMamba’s explainability, establishing it as a robust foundation for advancing single-cell transcriptomics in biological and biomedical research.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level (Korsunsky et al., 2019; Theodoris et al., 2023; Smith & Doe, 2024). This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing (Shen et al., 2023; Wen et al., 2023b; Nguyen & Tran, 2024). The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery (Gu & Dao, 2023; Yang et al., 2024; Li & Wang, 2024).
@@
Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise (Du et al., 2019; Zhao et al., 2023; Brown & Davis, 2024). Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data. Transformer-based architectures, such as scBERT (Yang et al., 2022) and scGPT (Cui et al., 2024a), have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis (Bian et al., 2024; Wen et al., 2023a; Garcia & Thompson, 2024).
@@
However, transformers exhibit critical limitations when applied to scRNA-seq data. The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes (Vaswani, 2017; Choromanski et al., 2020; Smith & Doe, 2024). Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions (Dao et al., 2022; Chen et al., 2023). These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences (Gu & Dao, 2023; Dao & Gu, 2024; Martinez & Hernandez, 2024).
@@
In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data. SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead (Shen et al., 2022; Gong et al., 2024a; Lee & Kim, 2024). Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness (Liu et al., 2024; Liang et al., 2024). We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis. Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models (Hao et al., 2024; Sun et al., 2024; Johnson & Lee, 2024). Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix D), highlighting its advantages in real-world biological applications.
@@ 2.1 Discretization of Gene Expression
Bin-based discretization, as used by scBERT (Yang et al., 2022), scGPT (Cui et al., 2024a), and scMulan (Bian et al., 2024), groups expression values into predefined bins. This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences. Additionally, binning can be sensitive to parameter selection, affecting downstream results.
Value projection(Szałata et al., 2024), adopted by scFoundation (Hao et al., 2024) and its backbone model xTrimoGene (Gong et al., 2024a), projects gene expression values into continuous embeddings rather than discrete categories. This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings. However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
Rank-based discretization, utilized by Geneformer (Theodoris et al., 2023), GeneCompass (Yang et al., 2024), tGPT(Shen et al., 2023) and LangCell (Zhao et al., 2024b), transforms gene expression values into ordinal rankings. This approach effectively captures relative expression levels and is more robust to batch effects and noise. Our method is based on rank-based discretization, as employed in Geneformer, which aligns naturally with biological processes such as regulatory interactions.
@@ 2.2 Model Architectures for Single-Cell Analysis
Transformer-based architectures, used by scBERT, scGPT, Geneformer, and scFoundation, have been successfully applied to single-cell data, leveraging the power of deep learning to model complex gene expression patterns. These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis. However, transformer(Vaswani, 2017) has inherent limitations. The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data. This issue arises because the self-attention mechanism used in transformer requires computing attention scores for all pairs of tokens, leading to inefficiencies in handling long sequences. Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions. Mamba(Gu & Dao, 2023) has been introduced to address the transformer efficiency problem. Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extraction(Sun et al., 2024), sequential recommendation(Liu et al., 2024), and time series forecasting(Liang et al., 2024). Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer. By leveraging state-space dynamics, Bi-Mamba can capture long-range dependencies effectively, making it well-suited for modeling gene regulatory interactions and cell state transitions. Furthermore, Bi-Mamba’s bidirectional processing enables the simultaneous consideration of upstream and downstream contexts, enhancing its ability to model complex dependencies in single-cell data. This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
More related work is provied in Appendix C.
@@ 3.1 Data Processing
$M_{ij}^{\text{norm}}=\frac{M_{ij}/\sum_{k=1}^{n}M_{ik}}{\text{t-digest}\{M_{kj}\mid M_{kj}>0\}}$
@@
$R_{i}=\operatorname{argsort}\left(-M_{ij}^{\text{norm}}\right),\quad\forall j$
@@ 3.2 Bi-Mamba Architecture
$z_{t}^{(l)}=\sigma(W^{(l)}[h_{t}^{(l)},\tilde{h}_{t}^{(l)}])$
$o_{t}^{(l)}=z_{t}^{(l)}\cdot h_{t}^{(l)}+(1-z_{t}^{(l)})\cdot\tilde{h}_{t}^{(l)}$
@@
$S^{(l+1)}=\text{Bi-Mamba}(S^{(l)}),\quad S^{(0)}=S$
@@ 3.3 Pretraining Objective
$\mathcal{L}_{\text{lang}}=-\frac{1}{M}\sum_{j=1}^{M}\log P(g_{j}\mid g_{1},g_{2},\dots,g_{j-1})$
@@
To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway (Oord et al., 2018; Sharma & Xu, 2023; Gundogdu et al., 2022). Pathways represent functional groupings of genes that collectively contribute to specific biological processes (Zien et al., 2000; García-Campos et al., 2015). Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar. Conversely, genes that belong to different pathways should be distinct in their embeddings. To achieve this, we define gene pairs within the same pathway as positive pairs (label $y_{ij}=1$) and gene pairs from different pathways as negative pairs (label $y_{ij}=0$).
@@
$\text{sim}(i,j)=\frac{\mathbf{h}_{i}\cdot\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$
@@
$\tilde{\text{sim}}(i,j)=\frac{\text{sim}(i,j)}{\tau}$
@@
$\mathcal{L}_{\text{pathway}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\tilde{\text{sim}}(i,j))}{\sum_{k\in\mathcal{N}(i)}\exp(\tilde{\text{sim}}(i,k))}$
@@
$\mathcal{L}=\mathcal{L}_{\text{lang}}+\gamma\mathcal{L}_{\text{pathway}}$
We set the $\gamma$ value to 0.1 in our experiments based on validation results from sample datasets.
@@ 4 Experiments
Our pretrianing dataset was constructed using single-cell RNA sequencing (scRNA-seq) data sourced from the CELLXGENE database, including raw count matrices and their corresponding metadata. The original dataset comprised 50,689,395 cells. To ensure data quality, duplicates were removed by retaining only entries where the feature is_primary_data was set to True. This step eliminated approximately 41% of the samples, resulting in 30,139,066 unique cells. Subsequently, the data was filtered to include 22,053 human protein-coding or miRNA genes. Cells expressing fewer than 200 genes were excluded to remove low-quality samples. The total gene expression counts for each cell were normalized to a fixed target value, and a logarithmic transformation ($log1p$) was applied to scale the data and reduce skewness. Following preprocessing, the dataset used for pretraining was finalized with 29,849,897 cells. For further refinement, a normalization factor was computed for each gene by calculating the non-zero median expression value of that gene across all cells. Genes were then ranked based on their normalized expression levels within each cell, and the top 2,048 or 4,096 gene indices were selected as input features for the pretraining stage.
@@ 4.1 Multi-batch Integration Task
The Avg-batch metric assesses the model’s ability to correct batch effects and integrate data across multiple batches, while the Avg-bio metric evaluates the preservation of biological variation and meaningful clustering of cell types after integration. As shown in Table 1 and Figure 3, batch correction and biological preservation often present a trade-off. Harmony, as a specialized method for batch effect correction, excels in eliminating batch effects but tends to ignore biological differences. In contrast, GeneMamba demonstrates superior performance by effectively mitigating batch effects while preserving biological information. This highlights its robust capability in multi-batch integration and its biological relevance for single-cell data analysis. For further details, please refer to Figures 9 and 10.
@@ 4.3 Gene Rank Reconstruction
Figure 4: Gene rank reconstruct results on PBMC12k dataset (a) Venn diagrams showing overlapping between input and output tokens in the pancreas dataset by three models: GeneMamba_U (unidirectional Mamba module as backbone), GeneFormer, GeneMamba (BiMamba module as backbone). (b) Density plots showing input and output ranking in the pancreas dataset by three models: GeneMamba_U, GeneFormer, GeneMamba.
The GeneMamba model demonstrates exceptional capability in reconstructing ranked gene orders, a critical requirement for single-cell analysis tasks where gene expression patterns reveal cellular states and transitions. The analysis begins by randomly selecting sample cells. From these cells, input gene tokens are extracted, and output tokens are generated using the GeneMamba model. Then the gene tokens are ranked based on their predicted likelihoods, with missing tokens filled by zeros to ensure sequence consistency. This approach preserves the structural integrity of the data and facilitates direct comparison between input and output rankings.
@@
Comparative analysis (Figure 4 from left to right) reveals the advantages of GeneMamba over other models: (1) The inclusion of the BiMamba module significantly enhances performance compared to the unidirectional module. This result demonstrates the necessity of bidirectional context extraction for effectively reconstructing ranked gene orders. (2) The GeneFormer model excels at predicting higher-ranked genes but struggles with lower-ranked genes, often treating them as noise due to their lower frequency. In contrast, GeneMamba captures both higher- and lower-ranked genes effectively, showcasing its robustness and sensitivity to varying gene expression levels. The metric results are summarized in Table 3. Additional results are provided in the Appendix Table 7. By reliably reconstructing ranked gene orders, GeneMamba offers a powerful solution for understanding complex gene expression patterns. Its ability to maintain rank integrity across varying data scales positions it as an invaluable tool for single-cell genomics.
4.4 Gene Correlation Analysis
Differentiating Positive and Negative Gene Pairs In the first experiment, GeneMamba’s ability to differentiate between ”positive” and ”negative” gene pairs was assessed using Gene2Vec embeddings. Nearly 30,000 gene pairs, evenly distributed as positive (1) or negative (0), were analyzed for similarity scores using cosine similarity and Pearson correlation. GeneMamba demonstrated a significant separation between the mean similarity scores of positive and negative pairs (Figure 5 a, b). Compared to baseline models, GeneMamba achieved a clearer distinction, emphasizing its capacity to accurately represent gene relationships.
Embedding Alignment To quantify discrimination power, we compared GeneMamba’s embeddings with those of scGPT and scFoundation. The ”distance” between positive and negative pair distributions was measured using Euclidean distance, KL divergence, and JS divergence. The result is shown in Appendix Figure 12. GeneMamba consistently outperformed the other models across all metrics, highlighting its superior ability to distinguish between gene pair types and reinforcing its robustness in modeling gene relationships.
Gene-Gene Topology Analysis The ability to capture gene-gene topologies was analyzed using adjacency matrices derived from embeddings. For the PBMC12K dataset, distances between embeddings of randomly selected gene pairs (100 cells in our case) were calculated and binarized based on threshold value. Then we build the topology graph based on the adjacent matrix (Figure  5c) for each evaluated model. Notably, while identifying unique model-specific topologies, all three models captured biologically significant genes and their interactions, such as RALA, DECR1, and CBL1. This suggests that the foundation models share a similar representation space.

### v2

# GeneMamba: An Efficient and Effective Foundation Model on Single Cell Data

@@ Abstract
Transformer-based architectures have shown promise across various domains but struggle with computational inefficiency and scalability. To address these challenges, we introduce GeneMamba, a novel model specifically designed for single-cell data analysis. GeneMamba incorporates the BiMamba module to efficiently capture gene context information and employs biologically informed loss functions during training. The model enables scalable processing of over 50 million cells while significantly reducing computational costs. It delivers strong performance in multi-batch integration, cell type annotation, and gene pair correlation analysis. Furthermore, reconstruction experiments highlight GeneMamba’s explainability, establishing it as a robust foundation for advancing single-cell transcriptomics in biological and biomedical research. By unifying bidirectional state space dynamics with rank-based gene tokenization, GeneMamba achieves both contextual depth and biological interpretability. These capabilities position GeneMamba as a scalable and principled backbone for future foundation models in single-cell omics.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level korsunsky2019fast; theodoris2023transfer. This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing shen2023generative; wen2023single; deconinck2021recent. The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery gu2023mamba; yang2024genecompass; li2024single.
@@
Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise du2019gene2vec; zhao2023large. Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data. Transformer-based architectures, such as scBERT yang2022scbert and scGPT cui2024scgpt, have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis bian2024scmulan; wen2023cellplm; szalata2024transformers.
@@
However, transformers exhibit critical limitations when applied to scRNA-seq data. The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes vaswani2017attention; choromanski2020rethinking. Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions dao2022flashattention; chen2023transformer. These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences gu2023mamba; mamba2.
@@
In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data. SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead shen2022universal; gong2024xtrimogene; ryu2023integration. Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness liu2024bidirectional; liang2024bi. We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis. Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models hao2024large; sun2024bidirectional; petegrosso2020machine. Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix F), highlighting its advantages in real-world biological applications.
@@ 2.1 Discretization of Gene Expression
Bin-based discretization, as used by scBERT yang2022scbert, scGPT cui2024scgpt, and scMulan bian2024scmulan, groups expression values into predefined bins. This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences. Additionally, binning can be sensitive to parameter selection, affecting downstream results.
Value projectionszalata2024transformers, adopted by scFoundation hao2024large and its backbone model xTrimoGene gong2024xtrimogene, projects gene expression values into continuous embeddings rather than discrete categories. This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings. However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
Rank-based discretization, utilized by Geneformer theodoris2023transfer, GeneCompass yang2024genecompass, tGPTshen2023generative and LangCell zhao2024langcell, transforms gene expression values into ordinal rankings. This approach effectively captures relative expression levels and is more robust to batch effects and noise. Our method is based on rank-based discretization, as employed in Geneformer, which aligns naturally with biological processes such as regulatory interactions.
@@ 2.2 Model Architectures for Single-Cell Analysis
Transformer-based architectures, used by scBERT, scGPT, Geneformer, and scFoundation, have been successfully applied to single-cell data, leveraging the power of deep learning to model complex gene expression patterns. These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis. However, transformervaswani2017attention has inherent limitations. The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data. This issue arises because the self-attention mechanism used in transformer requires computing attention scores for all pairs of tokens, leading to inefficiencies in handling long sequences. Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions. Mambagu2023mamba has been introduced to address the transformer efficiency problem. Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionsun2024bidirectional, sequential recommendationliu2024bidirectional, and time series forecastingliang2024bi. Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer. By leveraging state-space dynamics, Bi-Mamba can capture long-range dependencies effectively, making it well-suited for modeling gene regulatory interactions and cell state transitions. Furthermore, Bi-Mamba’s bidirectional processing enables the simultaneous consideration of upstream and downstream contexts, enhancing its ability to model complex dependencies in single-cell data. This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
More related work is provided in Appendix E.
@@ 3.1 Data Processing
$M_{ij}^{\text{norm}}=\frac{M_{ij}/\sum_{k=1}^{n}M_{ik}}{\text{t-digest}\{M_{kj}\mid M_{kj}>0\}}$ (1)
@@
$R_{i}=\operatorname{argsort}\left(-M_{ij}^{\text{norm}}\right),\quad\forall j$ (2)
@@ 3.2 Bi-Mamba Architecture
$z_{t}^{(l)}=\sigma(W^{(l)}[h_{t}^{(l)},\tilde{h}_{t}^{(l)}])$ (8)
$o_{t}^{(l)}=z_{t}^{(l)}\cdot h_{t}^{(l)}+(1-z_{t}^{(l)})\cdot\tilde{h}_{t}^{(l)}$ (9)
@@
$S^{(l+1)}=\text{Bi-Mamba}(S^{(l)}),\quad S^{(0)}=S$ (10)
@@ 3.3 Pretraining Objective
$\mathcal{L}_{\text{lang}}=-\frac{1}{M}\sum_{j=1}^{M}\log P(g_{j}\mid g_{1},g_{2},\dots,g_{j-1})$ (11)
@@
To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway oord2018representation; sharma2023regenne; gundogdu2022integrating. Pathways represent functional groupings of genes that collectively contribute to specific biological processes zien2000analysis; garcia2015pathway. Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar. Conversely, genes that belong to different pathways should be distinct in their embeddings. To achieve this, we define gene pairs within the same pathway as positive pairs (label $y_{ij}=1$) and gene pairs from different pathways as negative pairs (label $y_{ij}=0$).
@@
$\text{sim}(i,j)=\frac{\mathbf{h}_{i}\cdot\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$ (12)
@@
$\tilde{\text{sim}}(i,j)=\frac{\text{sim}(i,j)}{\tau}$ (13)
@@
$\mathcal{L}_{\text{pathway}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\tilde{\text{sim}}(i,j))}{\sum_{k\in\mathcal{N}(i)}\exp(\tilde{\text{sim}}(i,k))}$ (14)
@@
$\mathcal{L}=\mathcal{L}_{\text{lang}}+\gamma\mathcal{L}_{\text{pathway}}$ (15)
@@ 4 Experiments
Our pretrianing dataset was constructed using single-cell RNA sequencing (scRNA-seq) data sourced from the CELLXGENE database, including raw count matrices and their corresponding metadata. The original dataset comprised 50,689,395 cells. To ensure data quality, duplicates were removed by retaining only entries where the feature is_primary_data was set to True. This step eliminated approximately 41% of the samples, resulting in 30,139,066 unique cells. Subsequently, the data was filtered to include 22,053 human protein-coding or miRNA genes. Cells expressing fewer than 200 genes were excluded to remove low-quality samples. The total gene expression counts for each cell were normalized to a fixed target value, and a logarithmic transformation ($log1p$) was applied to scale the data and reduce skewness. Following preprocessing, the dataset used for pretraining was finalized with 29,849,897 cells. For further refinement, a normalization factor was computed for each gene by calculating the non-zero median expression value of that gene across all cells. Genes were then ranked based on their normalized expression levels within each cell, and the top 2,048 or 4,096 gene indices were selected as input features for the pretraining stage. The $\gamma$ value is set to 0.1 in our experiments based on validation results from sample datasets. The code and datasets are available at an anonymous address: https://anonymous.4open.science/r/GeneMamba-747D.
@@ 4.1 Multi-batch Integration Task
The Avg-batch metric assesses the model’s ability to correct batch effects and integrate data across multiple batches, while the Avg-bio metric evaluates the preservation of biological variation and meaningful clustering of cell types after integration. As shown in Table 1 and Figure 3, batch correction and biological preservation often present a trade-off. Harmony, as a specialized method for batch effect correction, excels in eliminating batch effects but tends to ignore biological differences. In contrast, GeneMamba demonstrates superior performance by effectively mitigating batch effects while preserving biological information. This highlights its robust capability in multi-batch integration and its biological relevance for single-cell data analysis. For further details, please refer to Figures 12 and 13.
@@ 4.3 Gene Rank Reconstruction
Figure 4: Gene rank reconstruct results on PBMC12k dataset. (a) Venn diagrams showing overlapping between input and output tokens in the pancreas dataset by three models: GeneMamba_U (unidirectional Mamba module as backbone), GeneFormer, GeneMamba (BiMamba module as backbone). (b) Density plots showing input and output ranking in the pancreas dataset by three models: GeneMamba_U, GeneFormer, GeneMamba.
The GeneMamba model exhibits strong capability in reconstructing ranked gene orders, a crucial requirement for single-cell analysis tasks where gene expression patterns reflect cellular states and transitions. This challenge has been highlighted in prior studies kedzierska2023assessing; boiarsky2023deep, and here we demonstrate the effectiveness of GeneMamba in addressing it. The analysis begins with the random selection of sample cells, from which input gene tokens are extracted. GeneMamba then generates output tokens, which are ranked according to their predicted likelihoods. To maintain sequence consistency, missing tokens are imputed with zeros. This strategy preserves the structural integrity of the data and enables a direct comparison between the input and output rankings.
@@
Comparative analysis (Figure 4 from left to right) reveals the advantages of GeneMamba over other models: (1) The inclusion of the BiMamba module significantly enhances performance compared to the unidirectional module. This result demonstrates the necessity of bidirectional context extraction for effectively reconstructing ranked gene orders. (2) The GeneFormer model excels at predicting higher-ranked genes but struggles with lower-ranked genes, often treating them as noise due to their lower frequency. In contrast, GeneMamba captures both higher- and lower-ranked genes effectively, showcasing its robustness and sensitivity to varying gene expression levels. The metric results are summarized in Table 3. Additional results are provided in the Appendix Table 7.